OpenStudy (anonymous):

f(x)=2x^2-4x+5 find f(-3)

OpenStudy (anonymous):

help xD

OpenStudy (lovelovelovehelp):

trying to figure it out first, sorry x/

OpenStudy (lovelovelovehelp):

It would be a solid 3 not -3

OpenStudy (anonymous):

really?

OpenStudy (lovelovelovehelp):

vertex Pt(-1,3)

OpenStudy (lovelovelovehelp):

x=1+sqrt(-3/2) or x=1-sqrt(-3/2). or i'm just confused :/ sorry

OpenStudy (lovelovelovehelp):

Put y=0 in y=2x^2-4x+5 and solve for x. 2x^2-4x+5=y. The same equation reversed for my convenience please. 2x^2-4x+5= 0, as y needs to be put to 0 2(x^2-2x)+5=0, rearranged . 2[(x-1)^2 -1}+5=0, rearranged. 2(x-1)^2-2 =-5. 2(x-1)^2=-5+2 . 2(x-1)^2=-3 (x-1)^2=-3/2 x-1= + or - sqrt(-3/2) x=1+sqrt(-3/2) or x=1-sqrt(-3/2).

Nnesha (nnesha):

f(-3) means x=-3 so substitute all x's for -3 and then solve that's it!

OpenStudy (lovelovelovehelp):

but it says find @Nnesha

Nnesha (nnesha):

yes correct find f(-3) so they are asking what would be the value of f(-3) when x=-3

OpenStudy (anonymous):

Can you show the work needed for the equation? @Nnesha

Nnesha (nnesha):

there is not much work for this here is an example \[\rm g(x)=2\color{ReD}{x}+4\] to find g(4) i would substitute x for 4 \[\rm g(4)=2\color{ReD}{(4)}+4\] and then just solve the right side